# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ A.1. Proof of Lemma 2 [Edge deletions that preserve $d$-Cayley-connectivity]
For any integer $n>0$, consider a disconnected compact subset $S\subset\mathbb{R}^{n}$ and let $S^{\prime}$ be one of its coordinate projections. If $S^{\prime}$ is connected, then there exists a point $x\in S^{\prime}$ whose fiber $S(x)$ is disconnected.
@@
Assume to the contrary that $S^{\prime}$ is connected and the fiber of any of its points through $S$ is connected. We can assume wlog that $S$ has exactly two connected components $S_{1}$ and $S_{2}$. Consider any two points $p\in S_{1}$ and $q\in S_{2}$ and let $p^{\prime}$ and $q^{\prime}$ be their projections in $S^{\prime}$. By assumption, there exists a path $P$ in $S^{\prime}$ between $p^{\prime}$ and $q^{\prime}$. Since $S_{1}$ and $S_{2}$ are compact, the fiber $S(x)$ of any point $x$ along $P$ is a subset of either $S_{1}$ or $S_{2}$. Hence, define $C(x)$ to be $1$ if $S(x)\subseteq S_{1}$, and $2$ otherwise. Also, let $Y$ be the set of all points $y$ along $P$ such that $C(y)=C(p^{\prime})$ and let $Z$ be the set of all points $z$ along $P$ such that $C(z)=C(q^{\prime})$. Note that $C(p^{\prime})\neq C(q^{\prime})$, and so $Y\cap Z=\emptyset$. However, this implies that $P$ is disconnected, which is a contradiction. Thus, the lemma is proved. ∎
Remark. Lemma 8 is true if we replace $S$ with any subset of a topological vector space.
@@
Assume to the contrary that $(G,F)$ and $(G\cup F,F^{\prime})$ are $d$-Cayley-connected but $(G,F\cup F^{\prime})$ is not. By definition, there exists a squared edge-length map $\ell$ such that the CCS $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ is disconnected. The CCS $\Omega^{d}_{F}(G,\ell)$ is connected, by assumption. Note that $\Omega^{d}_{F}(G,\ell)$ is a coordinate projection of $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$. Hence, letting $S=\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ and $S^{\prime}=\Omega^{d}_{F}(G,\ell)$, Lemma 8 shows that there exists a point $x\in S^{\prime}$ such that $S(x)$ is disconnected. However, this implies that there exists a squared edge-length map $\ell^{\prime}$ such that the CCS $\Omega^{d}_{F^{\prime}}(G\cup F,\ell^{\prime})$ is disconnected, which contradicts the fact that $(G\cup F,F^{\prime})$ is $d$-Cayley-connected. Therefore, $(G,F\cup F^{\prime})$ is $d$-Cayley-connected.
Finally, we show that if $(G,F\cup F^{\prime})$ is $d$-Cayley-connected, then so is $(G,F^{\prime})$. By assumption, the CCS $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ is connected for any squared edge-length map $\ell$. Since $\Omega^{d}_{F^{\prime}}(G,\ell)$ is a coordinate projection of $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$, it is connected. This completes the proof. ∎
A.2. Proof of Lemma 3 [Clique-gluing]
Lemma 3 is proved using Lemma 9, below. For any two induced subgraphs $H_{1}$ and $H_{2}$ of a graph $G$, $H_{1}\cap H_{2}$ is the subgraph of $G$ induced by $V(H_{1})\cap V(H_{2})$.
@@ A.2. Proof of Lemma 3 [Clique-gluing]
Consider a graph-nonedge pair $(G,f)$, a cs $C$ of $G\cup f$, and the set $\{G_{i}\}$ of all $C$-components of $G\cup f$ that contain $f$. Also, consider any linkage $(G,\ell)$ and let $\ell_{i}$ be the restriction of $\ell$ to the edges of $G_{i}\setminus f$. For any integer $d\geq 0$, the following equation is true:
@@
Let $I=\Omega^{d}_{f}(G,\ell)$ and $I_{i}=\Omega^{d}_{f}(G_{i}\setminus f,\ell_{i})$. First, we show that $I\subseteq\bigcap_{i}I_{i}$. For any point $x\in I$, there exists a $d$-realization of $(G,\ell)$ in which $f$ attains the squared length $x$. From this realization, we get a $d$-realization of each linkage $(G_{i}\setminus f,\ell_{i})$ in which $f$ attains the squared length $x$. Hence, $x$ is contained in $\bigcap_{i=1}^{n}I_{i}$.
Next, we show that $\bigcap_{i}I_{i}\subseteq I$. Let $\{H_{i}\}$ be the set of all $C$-components of $G\cup f$ and let $\ell^{\prime}_{i}$ be the restriction of $\ell$ to the edges of $H_{i}\setminus f$. Also, let $y$ be any point in $\bigcap_{i}I_{i}$ and let $p_{i}$ be any $d$-realization of $(H_{i}\setminus f,\ell^{\prime}_{i})\}$ such that if $H_{i}$ contains $f$, then $f$ attains the squared length $y$ in $p_{i}$. For any two $C$-components $H_{i}$ and $H_{j}$, observe that the restrictions of $p_{i}$ and $p_{j}$ to the vertex set of $H_{i}\cap H_{j}$ are equivalent up to translations, rotations, and overall reflection. Hence, we can construct a $d$-realization of $(G,\ell)$ in which $f$ attains the squared length $y$ by gluing all the realizations in $\{p_{i}\}$ together at their common intersections. Thus, $y$ is contained in $I$. ∎
An induced subgraph $H$ of a graph $G$ is preserved in a minor $[G]$ if each pair of its vertices is preserved in $[G]$. $H$ is weakly retained in $[G]$ if it is preserved and each of its nonedges is retained.
@@
For the other direction, let $\{H_{i}\}$ be the set of all $C$-components of $G\cup f$, $H^{\prime}_{i}=H_{i}\setminus f$, and $C^{\prime}=C\setminus f$. Assume wlog that $H_{1}$ contains $f$ and $(H_{1}\setminus,f)$ does not have the $d$-SIP. We will show that there exists an induced minor $[G]$ such that $([G],[f])$ does not have the $d$-SIP, and then apply the contrapositive of Lemma 1 to complete the proof. Let $[G]_{A}$ be the minor obtained by contracting all edges in $H^{\prime}_{i}\setminus C^{\prime}$ for each $H^{\prime}_{i}$ with $i\neq 1$. If $C^{\prime}$ does not contain $f$, then consider the minor $[G]_{B}$ obtained from $[G]_{A}$ by contracting any edge incident to each vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{A}$ with $i\neq 1$. Note that $[f]_{B}$ is a nonedge of $[G]_{B}$, $H^{\prime}_{1}$ is weakly retained in $[G]_{B}$, and $[G]_{B}=[H^{\prime}_{1}]_{B}$. Hence, since $(H^{\prime}_{1},f)$ does not have the $d$-SIP, neither does $([G]_{B},[f]_{B})$, as desired.
@@
Finally, if $C^{\prime}$ contains $f$, then consider the minor $[G]_{B}$ obtained from $[G]_{A}$ as follows. For any vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{A}$ adjacent to some vertex that is not an endpoint of $[f]_{A}$, contract the edge between two vertices. For any vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{A}$ that has degree $1$, contract the edge incident to it. Note that $[f]_{B}$ is a nonedge in $[G]_{B}$, $H^{\prime}_{1}$ is weakly retained in $[G]_{B}$, and $[G]_{B}$ can be obtained from $[H^{\prime}_{1}]_{B}$ by adding vertices of degree $2$ that neighbor the endpoints of $[f]_{B}$. Using these facts along with the triangle-inequality and the assumption that $(H_{1},f)$ does not have the $d$-SIP, it is easy to see that $([G]_{B},[f]_{B})$ does not have the $d$-SIP, as desired. This completes the proof. ∎
Appendix B Proofs of the converse direction propositions in Section 3
@@
B.1. Proof of Proposition 1 [Graphs with winged graph minors]
The proof of Proposition 1 requires Lemmas 10 and 11, below. The proof of Lemma 10 requires the following operation that transforms one minor $[G]_{A}$ of a graph $G$ into another minor $[G]_{B}$. Refer to Figure 16. Consider an edge $[uv]_{A}$ of $[G]_{A}$, the subgraph $H$ of $G$ induced by the set $[u]_{A}^{-1}$, a vertex $x\in[u]_{A}^{-1}$ with a neighbor in the set $[v]_{A}^{-1}$, and a connected component $J$ of $H\setminus x$. The minor $[G]_{B}$ obtained from $[G]_{A}$ by replacing $[u]_{A}^{-1}$ with $V(J)$ and $[v]_{A}^{-1}$ with $[v]_{A}^{-1}\cup(V(H)\setminus V(J))$ is said to be obtained via the vertex exchange from $[u]_{A}^{-1}$ to $[v]_{A}^{-1}$ that exchanges $x$ and fixes $J$.
Figure 16. A vertex exchange that exchanges $x$ and fixes $J$ is shown in red. A component exchange that exchanges $K$ is shown in purple. See the discussion below and Appendix B.2.
Figure 17. Graphs in Lemma 10.
Lemma 10 (Transforming a minor in Figure 17 into one in Figures 8 or 18).
Let $G$ be a $2$-connected graph and $e$ be one of its edges. If some minor $[G]_{A}$ is the graph in Figure 17(a) (resp. Figure 17(b)) such that $[e]_{A}$ is $w_{1}w_{2}$, then some minor $[G]_{B}$ is one of the graphs in Figures 8(a) (resp. Figure 8(b)) or 18(a)-18(d) (resp. 18(e)-18(i)) such that $[e]_{B}$ is $w_{1}w_{2}$. Furthermore, letting $\{w_{i}\}$ be the vertex sets of $[G]_{A}$ and $[G_{B}]$, we have $[w_{i}]_{A}^{-1}=[w_{i}]_{B}^{-1}$ for any $i\notin\{1,2\}$.
@@ B.1. Proof of Proposition 1 [Graphs with winged graph minors]
Let $e=uv$ and $[G]$ be a minor that is the graph in Figure 17(a) (resp. Figure 17(b)) and such that $[u]$ is $w_{1}$, $[v]$ is $w_{2}$, and $[u]^{-1}$ has minimum size. Let $H$ be the subgraph of $G$ induced by $[u]^{-1}$ and $x$ be any vertex in $[u]^{-1}\setminus\{u\}$ with a neighbor in $[v]^{-1}$, which exists since $G$ is $2$-connected. Also, let $J$ be the connected component of $H\setminus x$ that contains $u$. First, we show how to choose $x$ such that $J$ contains some vertex with a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$. If $J$ contains such a vertex, then our current choice for $x$ works. Otherwise, since there exists a path in $G\setminus x$ between $u$ and some vertex in $G\setminus([u]^{-1}\cup[v]^{-1})$, some connected component $K$ of $H\setminus x$ contains some vertex $y$ with a neighbor in $[v]^{-1}$ and some vertex with a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$. Note that the connected component of $H\setminus y$ that contains $u$ contains both $J$ and $x$. Hence, we can set $x$ to be $y$ and repeat this argument until $J$ has the desired property.
Next, consider the minor $[G]_{A}$ obtained via the vertex exchange from $[u]^{-1}$ to $[v]^{-1}$ that exchanges $x$ and fixes $J$. If $x$ does not have a neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$, then our above choice for $x$ ensures that $[G]_{A}$ is the graph in Figure 17(a) (resp. Figure 17(b)) such that $[u]_{A}$ is $w_{1}$ and $[v]_{A}$ is $w_{2}$. However, note that $[u]_{A}^{-1}$ is a proper subset of $[u]^{-1}$, which contradicts the minimality of $[u]^{-1}$. Therefore, $x$ has some neighbor in $G\setminus([u]^{-1}\cup[v]^{-1})$, and thus $[G]_{A}$ clearly has the desired properties. ∎
Figure 18. Graphs in Lemma 11. Solid line-segments are edges. Dashed line-segments are either edges or nonedges. Line-segments between $w_{2}$ and the blue circle represent an edge or nonedge between $w_{2}$ and some vertex in the circle.
For any edge $e$ of a graph $G$, if some minor $[G]$ is one of the graphs in Figure 18 such that $[e]$ is $w_{1}w_{2}$, then $G$ has an $e$-preserving forbidden minor.

### v2

@@ 4.1   Proofs of Tool Lemmas 2 and 3
For any dimension $n>0$, consider a disconnected compact subset $S\subset\mathbb{R}^{n}$ and let $S^{\prime}$ be one of its coordinate projections. If $S^{\prime}$ is connected, then there exists a point $x\in S^{\prime}$ whose fiber $S(x)$ is disconnected.
@@
Assume to the contrary that $S^{\prime}$ is connected and the fiber $S(x)$ of any point $x\in S^{\prime}$ is connected. We can assume wlog that $S$ has exactly two connected components $S_{1}$ and $S_{2}$. Consider any two points $p\in S_{1}$ and $q\in S_{2}$ and let $p^{\prime}$ and $q^{\prime}$ be their projections in $S^{\prime}$. By assumption, there exists a path $P$ in $S^{\prime}$ between $p^{\prime}$ and $q^{\prime}$. Since the fiber $S(x)$ of any point $x$ along $P$ is connected and the sets $S_{1}$ and $S_{2}$ are compact, $S(x)$ is a subset of either $S_{1}$ or $S_{2}$. Hence, define $C(x)$ to be $1$ if $S(x)\subseteq S_{1}$, and $2$ otherwise. Let $Y$ be the set of all points $y$ along $P$ such that $C(y)=C(p^{\prime})$ and let $Z$ be the set of all points $z$ along $P$ such that $C(z)=C(q^{\prime})$. By our assumptions, $C(p^{\prime})\neq C(q^{\prime})$ and $Y\cap Z=\emptyset$. However, this implies that $P$ is disconnected, which is a contradiction. Thus, the lemma is proved. ∎
Remark 2.
Lemma 7 is true if we replace $S$ with any subset of a topological vector space.
@@
Assume to the contrary that $(G,F)$ and $(G\cup F,F^{\prime})$ are $d$-Cayley-connected but $(G,F\cup F^{\prime})$ is not. By definition, there exists a squared edge-length map $\ell$ such that the CCS $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$ is disconnected. The CCS $\Omega^{d}_{F}(G,\ell)$, which is connected by assumption, is a coordinate projection of $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$. Hence, letting $S=\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$, $S^{\prime}=\Omega^{d}_{F}(G,\ell)$, and applying Lemma 7 shows that there exists a point $x\in S^{\prime}$ such that the fiber $S(x)$ is disconnected. However, this implies that there exists a squared edge-length map $\ell^{\prime}$ such that the CCS $\Omega^{d}_{F^{\prime}}(G\cup F,\ell^{\prime})$ is disconnected, which contradicts the fact that $(G\cup F,F^{\prime})$ is $d$-Cayley-connected. Therefore, $(G,F\cup F^{\prime})$ is $d$-Cayley-connected. The fact that $(G,F^{\prime})$ is also $d$-Cayley-connected follows from the observation that the CCS $\Omega^{d}_{F^{\prime}}(G,\ell)$ is a coordinate projection of $\Omega^{d}_{F\cup F^{\prime}}(G,\ell)$. This completes the proof. ∎
Lemma 3, about how $d$-SIP interacts with clique separators, is proved using Lemma 8, below. For any two induced subgraphs $H_{1}$ and $H_{2}$ of a graph $G$, $H_{1}\cap H_{2}$ is the subgraph of $G$ induced by $V(H_{1})\cap V(H_{2})$.
@@
Consider a graph-nonedge pair $(G,f)$, a clique separator $C$ of $G\cup f$, and the set $\{G_{i}\}$ of all $C$-components of $G\cup f$ that contain $f$. Also, consider any linkage $(G,\ell)$ and let $\ell_{i}$ be the restriction of $\ell$ to the edges of $G_{i}\setminus f$. For any dimension $d\geq 0$, the following equation holds:
@@
Let $I=\Omega^{d}_{f}(G,\ell)$ and $I_{i}=\Omega^{d}_{f}(G_{i}\setminus f,\ell_{i})$. The fact that $I\subseteq\bigcap_{i}I_{i}$ follows immediately from the definitions of these sets. To see that $\bigcap_{i}I_{i}\subseteq I$, let $\{H_{i}\}$ be the set of all $C$-components of $G\cup f$ and let $\ell^{\prime}_{i}$ be the restriction of $\ell$ to the edges of $H_{i}\setminus f$. Since $C$ is a clique, for any point in $\bigcap_{i}I_{i}$, corresponding $d$-realizations of each $(H_{i}\setminus f,\ell^{\prime}_{i})$ can be glued together on $C$ via translations, rotations, and overall reflection to obtain a $d$-realization of $(G,\ell)$. ∎
Recall from Section 3 that induced subgraph $H$ of a graph $G$ is preserved in a minor $[G]$ if each edge and nonedge of $H$ is preserved in $[G]$, and weakly retained in $[G]$ if $H$ is preserved in $[G]$ and each nonedge of $H$ is retained in $[G]$.
@@
For the other direction, let $\{H_{i}\}$ be the set of all $C$-components of $G\cup f$, $H^{\prime}_{i}=H_{i}\setminus f$, and $C^{\prime}=C\setminus f$. Assume wlog that $H_{1}$ contains $f$ and $(H_{1}\setminus,f)$ does not have the $d$-SIP. We will show that there exists an induced minor $[G]$ such that $([G],[f])$ does not have the $d$-SIP, and then apply the contrapositive of Lemma 1 to complete the proof. Let $[G]_{1}$ be the minor obtained by contracting all edges in $H^{\prime}_{i}\setminus C^{\prime}$ for each $H^{\prime}_{i}$ with $i\neq 1$. If $C^{\prime}$ does not contain $f$, then let the minor $[G]_{2}$ be obtained from $[G]_{1}$ by contracting any edge incident to each vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{1}$ with $i\neq 1$. Note that $[f]_{2}$ is a nonedge of $[G]_{2}$, $H^{\prime}_{1}$ is weakly retained in $[G]_{2}$, and $[G]_{2}=[H^{\prime}_{1}]_{2}$. Hence, since $(H^{\prime}_{1},f)$ does not have the $d$-SIP, neither does $([G]_{2},[f]_{2})$, as desired.
@@
Finally, if $C^{\prime}$ contains $f$, then consider the minor $[G]_{2}$ obtained from $[G]_{1}$ as follows. For any vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{1}$ adjacent to some vertex that is not an endpoint of $[f]_{1}$, contract the edge between these two vertices. For any vertex $[H^{\prime}_{i}\setminus C^{\prime}]_{1}$ that has degree $1$, contract the edge incident to it. Note that $[f]_{2}$ is a nonedge in $[G]_{2}$, $H^{\prime}_{1}$ is weakly retained in $[G]_{2}$, and $[G]_{2}$ can be obtained from $[H^{\prime}_{1}]_{2}$ by adding vertices of degree $2$ that neighbor the endpoints of $[f]_{2}$. Using these facts along with the triangle-inequality and the assumption that $(H_{1},f)$ does not have the $d$-SIP, it is easy to see that $([G]_{2},[f]_{2})$ does not have the $d$-SIP, as desired. This completes the proof. ∎
A clique separator $C$ of a graph $G$ is minimal, and referred to as a clique minimal separator (CMS) of $G$, if for any of its induced proper subgraphs $C^{\prime}$, the set of $C$-components of $G$ is not equal to the set of $C^{\prime}$-components of $G$.
Proof of Corollary 3.
Let $\mathcal{G}_{0}=\{G\cup f\}$ and $\mathcal{G}_{0},\dots,\mathcal{G}_{n}$ be a maximal sequence such that $\mathcal{G}_{i+1}$ is obtained from $\mathcal{G}_{i}$ by replacing one of its graphs $H$ that has a CMS $C$ with the $C$-components of $H$. It was shown in [1] that, for any such sequence, $\mathcal{G}_{n}$ is the set of all atoms of $G\cup f$. Lemma 3 shows that $(G,f)$ has the $d$-SIP if and only if, for each set $\mathcal{G}_{i}$ and each of its graphs $H$ that contain $f$, $(H\setminus f,f)$ has the $3$-SIP. This completes the proof. ∎
We are now ready to prove Proposition 2.
@@
Proposition 2 (Winged graph minors with singleton wing tips).
Consider a winged graph pair $(G,f)$ where $G\cup f$ is $2$-connected. If the converse direction of Theorem 4 is true for any graph-nonedge pair whose graph has strictly fewer vertices than $G$, then either $(G,f)$ has the $3$-SIP or some $f$-winged graph minor of $G\cup f$ has singleton wing tips.
@@
Let $f=uv$ and let $[G\cup f]$ be any $f$-winged graph minor with wing tips $s$ and $t$, as given by Lemma 6, which additionally minimizes the size of $[u]^{-1}\cup[v]^{-1}$. If $[u]^{-1}$ and $[v]^{-1}$ each contain exactly one vertex, then the lemma is proved. Otherwise, if $st$ is an edge of $G\cup f$, then $\{s,t\}$ is a CMS of $G\cup f$. Since $G\cup f$ has no $f$-preserving forbidden minor, by assumption, neither does any $\{s,t\}$-component $H$ of $G\cup f$ that contains $f$. Hence, since $|V(H)|<|V(G)|$, the converse direction of Theorem 4 along with Lemma 3 show that $(G,f)$ has the $3$-SIP. Therefore, assume that $st$ is a nonedge of $G\cup f$. We will show that either Lemma 2 can be used to prove that $(G,f)$ has the $3$-SIP or there exists an $f$-winged graph minor $[G\cup f]_{1}$ with wing tips such that $[u]_{1}^{-1}\cup[v]_{1}^{-1}$ is a proper subset of $[u]^{-1}\cup[v]^{-1}$, a contradiction. This will complete the proof.
Since $f$ is an edge of $G\cup f$, it is not $st$. A similar argument using the facts that $G\cup f$ has no $f$-preserving forbidden minor and $|V(H)|<|V(G)|$ along with the converse direction of Theorem 4 and Lemma 3 show that $(G\cup st,f)$ has the $3$-SIP. If $(G,st)$ has the $3$-SIP, then Lemma 2 states that $(G,f)$ has the $3$-SIP. Otherwise, note that $\{s,t\}$ is a clique separator of $G\cup st$, and so, by Lemma 3, there exists an $\{s,t\}$-component $H$ of $G\cup st$ that contains $st$ and such that $(H\setminus st,st)$ does not have the $3$-SIP. The fact that $|V(H)|<|V(G)|$ along with the contrapositive of the converse direction of Theorem 4 gives an $st$-preserving forbidden minor of $H$. Note that $H$ must contain $f$, or else $G\cup f$ clearly has an $f$-preserving forbidden minor, which is not the case. For the same reason, $H\cup f$ has no $f$-preserving forbidden minor. Also, since $G\cup f$ is $2$-connected, so is $H\cup f$. Therefore, Proposition 1 shows that $(H,f)$ is a winged graph pair with an $f$-winged graph minor $[H\cup f]_{2}$ whose vertex set is $\{[b_{i}]_{2}\}$.
Finally, since $H\cup f$ has an $st$-preserving forbidden minor, we can choose $[H\cup f]_{2}$ such that, wlog, $s$ is not contained in $[u]_{2}^{-1}\cup[v]_{2}^{-1}$. Assume that $s\in[b_{3}]_{2}^{-1}$. Using Lemma 6, we can additionally assume that $[H\cup f]_{2}$ has wing tips $s^{\prime}$ and $t^{\prime}$. Let $X=V(G\setminus H)$ and note that every vertex in $X$ is separated in $G$ from $V(H)\setminus\{s,t\}$ by $\{s,t\}$. Hence, since $st$ is an edge of $H\cup f$ and $s\in[b_{3}]_{2}^{-1}$, there exists an $f$-winged graph minor $[G\cup f]_{1}$ with wing tips $s^{\prime}$ and $t^{\prime}$ and whose vertex set is $\{[a_{i}]_{1}\}$, where $[a_{3}]_{1}^{-1}=[b_{3}]_{2}^{-1}\cup X$ and $[a_{i}]_{1}^{-1}=[b_{i}]_{2}^{-1}$ for any $i\neq 3$. Thus, since $[u]_{1}^{-1}\cup[v]_{1}^{-1}=[u]_{2}^{-1}\cup[v]_{2}^{-1}$ and $[u]_{2}^{-1}\cup[v]_{2}^{-1}\subsetneq V(H)\subseteq[u]^{-1}\cup[v]^{-1}$, we arrive at the desired contradiction. ∎
Next we move to Proposition 4, the last one in the first part of the proof of the converse direction of the main Theorem 4. For any graph $G$, subset $U\subseteq V(G)$, and vertices $s,t\in V(G)\setminus U$, $U$ is an $st$-separator of $G$ if no connected component of $G\setminus U$ contains both $s$ and $t$, and additionally minimal if no proper subset of $U$ is an $st$-separator of $G$. Let $f_{\star}=uv$. A pair of vertices $xy$ of $G_{\star}$ is a pin of $[G_{\star}\cup f_{\star}]$ if $\{x,y\}$ is a minimal $uv$-separator of $G_{\star}\setminus[f_{\star}]^{-1}$ such that either $x\in[a_{3}]^{-1}$ and $y\in[a_{4}]^{-1}$ or $x\in[a_{5}]^{-1}$ and $y\in[a_{6}]^{-1}$. Note that the paths in $G_{\star}\setminus[f_{\star}]^{-1}$ given by the existence of $[G_{\star}\cup f_{\star}]$ show that any minimal $uv$-separator of $G_{\star}\setminus[f_{\star}]^{-1}$ that is contained in $V(G_{\star})\setminus\{[u]^{-1},[v]^{-1}\}$ and has size at most two is a pin of $[G_{\star}\cup f_{\star}]$. Proposition 3, stated below and used to prove Proposition 4, shows that $[G_{\star}\cup f_{\star}]$ has pins $xy$ and $wz$ with $x\in[a_{3}]^{-1}$, $y\in[a_{4}]^{-1}$, $w\in[a_{5}]^{-1}$, and $z\in[a_{6}]^{-1}$. See Figure 12(c). Proposition 4 then allows us to assume that each pin of $[G_{\star}\cup f_{\star}]$ is an edge of $G_{\star}\cup f_{\star}$, in which case we call $[G_{\star}\cup f_{\star}]$ chained. See Figure 12(d).
Proposition 3 is proved using Lemma 9, below. The proof of this lemma involves inspecting many simple cases, and so it is given in A.1.
Lemma 9 (Inductive structure of winged graph minors).
Let $(G,f)$ be a winged graph pair, $[G\cup f]$ be an $f$-winged graph minor whose vertex set is $\{[c_{i}]\}$, and $X\subseteq[c_{3}]^{-1}$ and $Y\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{1}]^{-1}$. Also, let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{5}]^{-1}\cup[c_{6}]^{-1}$ if $[G\cup f]$ is the graph in Figure 8(a), otherwise; let $A\subseteq[c_{3}]^{-1}$ and $B\subseteq[c_{4}]^{-1}$ be the subsets of all vertices with neighbors in $[c_{5}]^{-1}\cup[c_{7}]^{-1}$ and $[c_{6}]^{-1}\cup[c_{7}]^{-1}$ if $[G\cup f]$ is the graph in Figure 8(b). If some set in $\{X,Y\}$ and some set in $\{A,B\}$ both have size at least two, then there exists an $f$-winged graph minor $[G\cup f]_{1}$ whose vertex set $\{[d_{i}]_{1}\}$ is such that $[d_{3}]_{1}^{-1}\cup[d_{4}]_{1}^{-1}$ is a proper subset of $[c_{3}]^{-1}\cup[c_{4}]^{-1}$ and both $[f]_{1}^{-1}$ and $[f]^{-1}$ contain the same set of edges of $G$.
Proposition 3 (Pins of a winged graph minor).
For any winged graph pair $(G,f)$, any $f$-winged graph minor $[G\cup f]$, whose vertex set is $\{[a_{i}]\}$, has pins $xy$ and $wz$ with $x\in[a_{3}]^{-1}$, $y\in[a_{4}]^{-1}$, $w\in[a_{5}]^{-1}$, and $z\in[a_{6}]^{-1}$.